DIRECT-POTENTIAL-FIT ANALYSIS FOR Li_{2}(a^{3}Σ_{u}^{+}) AND EXTENSIONS OF THE ‘MLR’ POTENTIAL FUNCTION MODEL

\underline{NIKESH S. DATTANI} and ROBERT J. LE ROY, Guelph-Waterloo Centre for Graduate Work in Chemistry and Biochemistry, University of Waterloo, Waterloo, Ontario N2L 3G1, Canada; AMANDA J. ROSS, Université Lyon 1; CNRS; LASIM UMR 5579, 43 Bd du 11 novembre 1918, F-69622 Villeurbanne, France; COLAN LINTON, Physics Department, University of New Brunswick, Fredericton, NB E3B 5A3, Canada.

While a number of studies of the weakly bound a^{3}Σ_{u}^{+} state of ^{6,6}Li_{2} and ^{7,7}Li_{2} have been reported, the only potential functions obtained for these systems were based on point-wise semiclassical RKR curves generated from Dunham or neardissociation expansions for the vibrational energies and B_{v} constants.^{a} Moreover, to date the data for ^{6,6}Li_{2} and ^{7,7}Li_{2} have always been treated independently, so the effect of Born-Oppenheimer breakdown in this system is unknown. The present work reports a combined-isotopologue direct-potential-fit analysis of all available fluorescence and PAS data for the 1^{3}Σ_{g}^{+} − a^{3}Σ_{u}^{+} and 2^{3}Π_{g} − a^{3}Σ_{u}^{+} systems of ^{6,6}Li_{2} and ^{7,7}Li_{2}. The analytic potential energy functions used to characterize the a^{3}Σ_{u}^{+} and 1^{3}Σ_{g}^{+} states are extended versions of the ‘Morse-Long-Range’ (MLR) potential model which explicitly incorporates the theoretically-known inverse-power long-range behaviour within a unified potential function form.^{b} ‘Adiabatic’ Born-Oppenheimer breakdown functions are required to yield a consistent analysis of the data for the two isotopologues, and they yield isotopologue-dependent well depths for these two states. The −C_{3}/r^{3} limiting long-range behaviour of the 1^{3}Σ_{g}^{+}-state potential presented a challenge to the use of the MLR potential function form with two or more long-range terms, and this led to better understanding and a significant extension of the MLR model.

a(a) F. Martin et al., Spectrochim. Acta 44A, 1369 (1988); (b) C. Linton et al., J. Chem. Phys. 91, 6036 (1989); (c) W.T. Zemke and W.C. Stwalley, J. Phys. Chem. 97, 2053 (1993); (d) A.J. Moerdijk et al., Phys. Rev. Lett. 72, 40 (1994); (e) C. Linton et al., J. Mol. Spectrosc. 196, 20 (1999).

bR.J. Le Roy, Y. Huang and C. Jary, J. Chem. Phys. 125, 164310 (2006); R.J. Le Roy and R.D.E. Henderson, Mol. Phys. 105, 663 (2007).